OpenStudy (anonymous):

How do I simplify: x-3 3-x ----- - ---- x^2-4 4-x^2

OpenStudy (phi):

make the denominators the same. one way is multiply the second fraction top and bottom by -1

OpenStudy (anonymous):

\[(x-3\div x ^{2}-4) - (3-x \div 4-x ^{2})\]

OpenStudy (anonymous):

that's what I did. And I got x-3 3+x ----- + ------- x^2-4 (x^2-4)

OpenStudy (phi):

it is still a minus between the two fractions.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

the top was (3-x) times -1 you get -1(3-x)= -3+x = x+3

OpenStudy (anonymous):

ohhh.

OpenStudy (phi):

Looks good.

OpenStudy (anonymous):

ohh, okay. thank you!

OpenStudy (phi):

Though, to not confuse myself, I would re-write the original problem as \[ \frac{x-3)}{x^2-4} + \frac{-(3-x)}{4-x^2} \]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

isn't it suppose to be subtraction?

OpenStudy (phi):

then multiply top and bottom by -1 to get \[\frac{(x-3)}{x^2-4} + \frac{(3-x)}{x^2-4}\]

OpenStudy (phi):

yes by x - y is the same as x + (-y)

OpenStudy (anonymous):

oh. okay. thank you!

OpenStudy (phi):

don't worry about it if it is not helpful

OpenStudy (phi):

Final answer is 0

OpenStudy (anonymous):

it is?

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

the answer is 6/x^2-4.. o.o

OpenStudy (anonymous):

on the back of my book.

OpenStudy (anonymous):

hello?

OpenStudy (phi):

Did you post same equation as in the book?

OpenStudy (anonymous):

yes.. x+3 3-x ------ - ----- x^2 - 4 4-x^2

OpenStudy (phi):

which is the same as \[ \frac{x+3}{x^2-4} + \frac{3-x}{x^2-4} \] \[ = \frac{x+3+3-x}{x^2-4} \]

OpenStudy (phi):

looks like your book is right!

OpenStudy (anonymous):

ohhhh!!

OpenStudy (phi):

I used the minus sign to change the denominator. But it should work if we multiply top and bottom by -1 and go from there.

OpenStudy (anonymous):

yup, I see it! :P Thank you!

OpenStudy (phi):

I think I don't like minus signs!

OpenStudy (phi):

OK, I see how you confused me. In the original problem you have (x-3) in the first fraction. Now you have (x+3).

OpenStudy (anonymous):

ohhh, yeah. Sorry!! D: